# Grounded in Context: Retrieval-Based Method for Hallucination Detection

### v1

@@ Abstract
Despite advancements in grounded content generation, production Large Language Models (LLMs) based applications still suffer from hallucinated answers. We present ”Grounded in Context” - Deepchecks’ hallucination detection framework, designed for production-scale long-context data and tailored to diverse use cases, including summarization, data extraction, and RAG. Inspired by RAG architecture, our method integrates retrieval and Natural Language Inference (NLI) models to predict factual consistency between premises and hypotheses using an encoder-based model with only a 512-token context window. Our framework identifies unsupported claims with an F1 score of 0.83 in RAGTruth’s response-level classification task, matching methods that trained on the dataset, and outperforming all comparable frameworks using similar-sized models.
@@ 1 Introduction
To address these challenges, we developed Grounded in Context, a RAG-inspired methodology for hallucination detection. Our approach decomposes output into factual statements and retrieves a dedicated context for each statement. Statement-context pairs are evaluated for factual consistency, with individual scores aggregated into a comprehensive metric. Our methodology demonstrates superior performance compared to models of comparable size when evaluated on the RAGTruth dataset [2], despite the exclusion of its training corpus from our method’s training data.

### v3

# ORION Grounded in Context: Retrieval-Based Method for Hallucination Detection

@@ Abstract
Despite advancements in grounded content generation, production Large Language Models (LLMs) based applications still suffer from hallucinated answers. We present ”Grounded in Context” - a member of Deepchecks’ ORION (Output Reasoning-based InspectiON) family of lightweight evaluation models. It is our framework for hallucination detection, designed for production-scale long-context data and tailored to diverse use cases, including summarization, data extraction, and RAG. Inspired by RAG architecture, our method integrates retrieval and Natural Language Inference (NLI) models to predict factual consistency between premises and hypotheses using an encoder-based model with only a 512-token context window. Our framework identifies unsupported claims with an F1 score of 0.83 in RAGTruth’s response-level classification task, matching methods that trained on the dataset, and outperforming all comparable frameworks using similar-sized models.
@@ 1 Introduction
To address these challenges, we present ORION (Output Reasoning-based InspectiON) - a family of lightweight evaluation models designed to assess the quality of LLM and SLM outputs. It performs structured, multistep analysis across dimensions such as factual consistency, information density, task relevance, and avoidance. ORION enables precise, reliable, and explainable evaluation pipelines critical for quality assurance, compliance, and iterative development of language-based systems.
A crucial member of the ORION family is ”Grounded in Context”, a RAG-inspired methodology for hallucination detection. Our approach decomposes outputs into factual statements and retrieves a dedicated context for each statement. Statement-context pairs are evaluated for factual consistency, with individual scores aggregated into a comprehensive metric. Our methodology demonstrates superior performance compared to models of comparable size when evaluated on the RAGTruth dataset [2], despite the exclusion of its training corpus from our method’s training data.